# Foundations of Adaptive High-Level Tight Control of Prostate Cancer: A Path from From Terminal Disease to Chronic Condition

### v2

@@ V.1 $\text{PSA}_{\text{thr}}/{\text{PSA}_{K}}_{\mu}\leq 1/\alpha_{RS}$
Here, we search for the (global) maximum of the total achievable time (with treatment) below the PSA threshold, $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$ over this multi-dimensional space $(\eta_{\uparrow},\eta_{\downarrow})$. We employ a Bayesian optimization approach [39] to efficiently navigate through the policy parameter space by data-driven extrapolation (kriging [40]) for the two-dimensional landscape $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$ for patient $\mu$ with a given $\text{PSA}_{\text{thr}}$. This method, when applied to an objective function to be extremize, begins by evaluating that function at a number of random positions so as to (approximately) gauge its global features (using a surrogate model e.g. a Gaussian process [41]). It then uses Bayesian inference to predict the optimal next function evaluation location, balancing exploitation and exploration, and evaluate the function there. This iterative step continuously updates its estimation of the function, keeps finding and checking new locations and corresponding values, until the difference between two consecutive locations becomes sufficiently small (for a practical resolution of the parameter space).
@@ V.1.1 Bayesian Optimization
using Bayesian optimization [39] with Gaussian process regression (GPR) [41]. Consider a search within the region $(\eta_{\uparrow},\eta_{\downarrow})\in[0,0.90]\times[0,0.90]$, where the resolution in each direction is $\delta\eta=0.01$. Note that the upper bounds for $\eta_{\uparrow}$ and $\eta_{\downarrow}$ are both set at $0.90$ instead of $1.00$, as smaller differences (less that $10\%$) are clinically difficult to monitor [21]. We start by choosing $n_{\text{eval}}=10$ random points
@@
Bayesian optimization utilize this GPR extrapolation [^4] We utilize the function fitrgp in MatLab R2023a [29] for this task, using the ardmatern32 kernel function and the “exact fitting” option. to identify the next candidate points in the search space, $(\eta_{\uparrow}^{(n_{\text{eval}}+1)},\eta_{\downarrow}^{(n_{\text{eval}}+1)})$, where the total achievable time is likely to be maximized with some confidence. This is done by balancing exploration (searching in regions with high uncertainty) and exploitation (focusing on regions predicted to have high values):
@@ V.1.2 Optimal Treatment for Maximum $\tau_{\text{tot}}$
  • Panel (A,D): Cancer progression under continuous Abiraterone treatment often results in resistant cells rapidly overtaking the population, causing the cancer to exceed the thresholds faster. Note that continuous therapy is a special case of adaptive therapy in which the off-drug policy parameter $\eta_{\downarrow}$ is set to 00.
  • Panel (B,E): Cancer progression under a (non-optimal) adaptive Abiraterone treatment, in which $\text{PSA}_{\uparrow}$ is as high as possible and $\text{PSA}_{\downarrow}$ is set to $50\%$ of $\text{PSA}_{\uparrow}$. This treatment has been used in studies and clinical trials [22]. Since the drug is not used to suppress sensitive cells constantly, these cells are allowed to recover and keep competing with the resistant ones longer, hence the total achievable time is extended.
@@ V.2 $\text{PSA}_{\text{thr}}/{\text{PSA}_{K}}_{\mu}>1/\alpha_{RS}$
What we have observed in this Section and in Section V.1.2 is that HLTC treatment typically leads to improved outcomes, either by maximizing the total drug response time for low $\text{PSA}_{\text{thr}}$ or by transforming the disease into a chronic condition for sufficiently high $\text{PSA}_{\text{thr}}$. This simple insight could serve as a valuable guide for enhancing current adaptive chemotherapy treatments. We suggest that clinicians might explore the possibility of implementing HLTC, i.e. waiting for the disease to advance further rather than initiating treatment too early. It is worth mentioning that this kind of “watchful waiting” approach before treatment is not uncommon in practice [1], at least historically. Furthermore, even without employing the “watchful waiting” approach (i.e., disregarding the high level component of HLTC), maintaining the current method for determining the appropriate PSA level for the initial drug treatment (that is, $\text{PSA}_{\uparrow}$) but implementing a tighter control, may also change prostate cancer into a chronic disease for many patients (our simulation indicates that, with this approach, 19 out of 32 patients could potentially become chronic).
@@ VI Discussion
Cancer has posed an immense challenge to modern medicine due to its complex nature and the absence of a definitive cure [42], casting a shadow of devastation upon those affected. This disease can wreak havoc on the human body, resulting in a multitude of severe impacts before leading to the eventual demise. While prostate cancer treatments can prolong life-expectancy, they can also be extremely aggressive and cause debilitating side effects [43]. Chemotherapy, a common treatment, frequently induces hair loss, nausea, and extreme fatigue, significantly impacting not only the physical but also the emotional well-being of patients [44]. It is crucial to understand that augmented drug administration does not always result in an increased lifespan, even at the cost of more suffering. Paradoxically, it appears plausible to significantly extend the patient life while concurrently reducing the treatment ratio (the proportion of time a drug is given compared to the total treatment duration) throughout the treatment course via adaptive chemotherapy [20]. In other words, a longer and better life may be achievable.
@@
In this work, we use a systematic statistical method, LBEB approach for NLME, to estimate patient-informed parameters for a Stackelberg game-theoretic multi-population model that quantitatively describes prostate cancer progression [22]. We obtain estimates for common parameters across all patients as well as the probability distribution for patient-specific parameters in Section IV. We then employ a Bayesian Optimization approach to identify the optimal adaptive chemotherapeutic treatment policy for a single drug (Abiraterone) aimed at maximizing the time before the patient begins to show symptoms, under the assumption that we know how advanced the cancer needs to be for symptoms to become manifest, in Section V. We show that a high level tight control (HLTC) treatment, in which the trigger signals (i.e., the biomarker levels) for drug administration and cessation are both high and close together, typically yields the best outcomes, as demonstrated through both computer-assisted and theoretical means (see Appendix B). Furthermore, we demonstrate that it may be possible to transform prostate cancer from a terminal disease into a chronic condition for most patients in Section V.2.
@@
Following this work, there are many directions for further exploration. First and foremost is realizing the potential benefits of HLTC adaptive therapeutic treatments in laboratory studies and clinical practice. Work along this direction is already in progress. On the numerical investigation front, we have only optimize (via Bayesian Optimization) a very simple treatment policy, in which there are only two states to be adjusted in drug administration: either no drug or the maximum dosage rate. Allowing for more structured, possibly continuous in time – not just on/off – as proposed in [45], will provide a much richer protocol space to optimize in, and hopefully, better solution to discover. Additional factors can also be considered, such as the extreme toxicity of Docetaxel [20]. Too frequent dosages can be intolerable to the patient’s body. Thus, for the treatment policy, limits should be imposed on the administration of Docetaxel within a specific timeframe to ensure patient safety, which is indeed often the case in clinical practices.
@@
Better analysis on more realistic models of cancer could also lead to new insights and novel discoveries regarding collective behavior emergence from population dynamics under ever-changing conditions [46]. Stochastic therapies (in the spirit of the Kapitza stochastic stabilization of the inverted pendulum [47, 48]) might prove more effective in targeting cancerous cells compared to periodic ones (a phenomenon observed in a robotic model system [49]). Importantly, we know that spatial dynamics are crucial in evolutionary systems [50]. Considering the metastatic nature of cancer cells, capturing how cells migrate (and also the topology of the environment the cells live in [51, 52]) in the mathematical models should be of critical importance as we strive for modeling fidelity.
@@ B.1 Assumptions and Approximations
[…]
@@
The right hand side is a finite positive value very close to 00, which is also increasing with $\epsilon$. This means that the tighter the control (i.e. the smaller the control window), the more difficult it becomes to maintain stabilization Eq. (30), as it reduces the allowable range for $x_{S}$. Therefore, when the adaptive treatment starts failing to suppress the cancer growth (i.e. the inequality does not hold anymore), we can make the following approximation:

### v3

@@ V.1 $\text{PSA}_{\text{thr}}/{\text{PSA}_{K}}_{\mu}\leq 1/\alpha_{RS}$
Here, we search for the (global) maximum of the total achievable time (with treatment) below the PSA threshold, $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$ over this multi-dimensional space $(\eta_{\uparrow},\eta_{\downarrow})$. We employ a Bayesian optimization approach [frazier2018tutorial] to efficiently navigate through the policy parameter space by data-driven extrapolation (kriging [oliver1990kriging]) for the two-dimensional landscape $\tau_{\text{tot}}^{\{\mu\},\text{PSA}_{\text{thr}}}(\eta_{\uparrow},\eta_{\downarrow})$ for patient $\mu$ with a given $\text{PSA}_{\text{thr}}$. This method, when applied to an objective function to be extremize, begins by evaluating that function at a number of random positions so as to (approximately) gauge its global features (using a surrogate model e.g. a Gaussian process [schulz2018tutorial]). It then uses Bayesian inference to predict the optimal next function evaluation location, balancing exploitation and exploration, and evaluate the function there. This iterative step continuously updates its estimation of the function, keeps finding and checking new locations and corresponding values, until the difference between two consecutive locations becomes sufficiently small (for a practical resolution of the parameter space).
@@ V.1.1 Bayesian Optimization
using Bayesian optimization [frazier2018tutorial] with Gaussian process regression (GPR) [schulz2018tutorial]. Consider a search within the region $(\eta_{\uparrow},\eta_{\downarrow})\in[0,0.90]\times[0,0.90]$, where the resolution in each direction is $\delta\eta=0.01$. Note that the upper bounds for $\eta_{\uparrow}$ and $\eta_{\downarrow}$ are both set at $0.90$ instead of $1.00$, as smaller differences (less that $10\%$) are clinically difficult to monitor [hansen-read]. We start by choosing $n_{\text{eval}}=10$ random points
@@
Bayesian optimization utilize this GPR extrapolation [^5] We utilize the function fitrgp in MatLab R2023a [MATLAB] for this task, using the ardmatern32 kernel function and the “exact fitting” option. to identify the next candidate points in the search space, $(\eta_{\uparrow}^{(n_{\text{eval}}+1)},\eta_{\downarrow}^{(n_{\text{eval}}+1)})$, where the total achievable time is likely to be maximized with some confidence. This is done by balancing exploration (searching in regions with high uncertainty) and exploitation (focusing on regions predicted to have high values):
@@ V.1.2 Optimal Treatment for Maximum $\tau_{\text{tot}}$
  • Panel (A,D): Cancer progression under continuous Abiraterone treatment often results in resistant cells rapidly overtaking the population, causing the cancer to exceed the thresholds faster. Note that continuous therapy is a special case of adaptive therapy in which the off-drug policy parameter $\eta_{\downarrow}$ is set to 0.
  • Panel (B,E): Cancer progression under a (non-optimal) adaptive Abiraterone treatment, in which $\text{PSA}_{\uparrow}$ is as high as possible and $\text{PSA}_{\downarrow}$ is set to $50\%$ of $\text{PSA}_{\uparrow}$. This treatment has been used in studies and clinical trials [zhang2022evolution]. Since the drug is not used to suppress sensitive cells constantly, these cells are allowed to recover and keep competing with the resistant ones longer, hence the total achievable time is extended.
@@ V.2 $\text{PSA}_{\text{thr}}/{\text{PSA}_{K}}_{\mu}>1/\alpha_{RS}$
What we have observed in this Section and in Section V.1.2 is that HLTC treatment typically leads to improved outcomes, either by maximizing the total drug response time for low $\text{PSA}_{\text{thr}}$ or by transforming the disease into a chronic condition for sufficiently high $\text{PSA}_{\text{thr}}$. This simple insight could serve as a valuable guide for enhancing current adaptive chemotherapy treatments. We suggest that clinicians might explore the possibility of implementing HLTC, i.e. waiting for the disease to advance further rather than initiating treatment too early. It is worth mentioning that this kind of “watchful waiting” approach before treatment is not uncommon in practice [chapple2002watchful], at least historically. Furthermore, even without employing the “watchful waiting” approach (i.e., disregarding the high level component of HLTC), maintaining the current method for determining the appropriate PSA level for the initial drug treatment (that is, $\text{PSA}_{\uparrow}$) but implementing a tighter control, may also change prostate cancer into a chronic disease for many patients (our simulation indicates that, with this approach, 19 out of 32 patients could potentially become chronic).
@@ VI Discussion
Cancer has posed an immense challenge to modern medicine due to its complex nature and the absence of a definitive cure [hayden2008cancer], casting a shadow of devastation upon those affected. This disease can wreak havoc on the human body, resulting in a multitude of severe impacts before leading to the eventual demise. While prostate cancer treatments can prolong life-expectancy, they can also be extremely aggressive and cause debilitating side effects [schirrmacher2019chemotherapy]. Chemotherapy, a common treatment, frequently induces hair loss, nausea, and extreme fatigue, significantly impacting not only the physical but also the emotional well-being of patients [aslam2014side]. It is crucial to understand that augmented drug administration does not always result in an increased lifespan, even at the cost of more suffering. Paradoxically, it appears plausible to significantly extend the patient life while concurrently reducing the treatment ratio (the proportion of time a drug is given compared to the total treatment duration) throughout the treatment course via adaptive chemotherapy [west2019multidrug]. In other words, a longer and better life may be achievable.
@@
In this work, we use a systematic statistical method, LBEB approach for NLME, to estimate patient-informed parameters for a Stackelberg game-theoretic multi-population model that quantitatively describes prostate cancer progression [zhang2022evolution]. We obtain estimates for common parameters across all patients as well as the probability distribution for patient-specific parameters in Section IV. We then employ a Bayesian Optimization approach to identify the optimal adaptive chemotherapeutic treatment policy for a single drug (Abiraterone) aimed at maximizing the time before the patient begins to show symptoms, under the assumption that we know how advanced the cancer needs to be for symptoms to become manifest, in Section V. We show that a high level tight control (HLTC) treatment, in which the trigger signals (i.e., the biomarker levels) for drug administration and cessation are both high and close together, typically yields the best outcomes, as demonstrated through both computer-assisted and theoretical means (see Appendix B). Furthermore, we demonstrate that it may be possible to transform prostate cancer from a terminal disease into a chronic condition for most patients in Section V.2.
@@
Following this work, there are many directions for further exploration. First and foremost is realizing the potential benefits of HLTC adaptive therapeutic treatments in laboratory studies and clinical practice. Work along this direction is already in progress. On the numerical investigation front, we have only optimize (via Bayesian Optimization) a very simple treatment policy, in which there are only two states to be adjusted in drug administration: either no drug or the maximum dosage rate. Allowing for more structured, possibly continuous in time – not just on/off – as proposed in [cunningham2018optimal], will provide a much richer protocol space to optimize in, and hopefully, better solution to discover. Additional factors can also be considered, such as the extreme toxicity of Docetaxel [west2019multidrug]. Too frequent dosages can be intolerable to the patient’s body. Thus, for the treatment policy, limits should be imposed on the administration of Docetaxel within a specific timeframe to ensure patient safety, which is indeed often the case in clinical practices.
@@
Better analysis on more realistic models of cancer could also lead to new insights and novel discoveries regarding collective behavior emergence from population dynamics under ever-changing conditions [phan2021doesn]. Stochastic therapies (in the spirit of the Kapitza stochastic stabilization of the inverted pendulum [kapitza1951dynamic, sanz2008stabilizing]) might prove more effective in targeting cancerous cells compared to periodic ones (a phenomenon observed in a robotic model system [wang2022robots]). Importantly, we know that spatial dynamics are crucial in evolutionary systems [wu2013cell]. Considering the metastatic nature of cancer cells, capturing how cells migrate (and also the topology of the environment the cells live in [phan2020bacterial, phan2024social]) in the mathematical models should be of critical importance as we strive for modeling fidelity.
VII Acknowledgement
This work was supported by the US National Science Foundation (PHY-1659940 and PHY-1734030), and the Johns Hopkins University Discovery Award 2023-2024.
@@ B.1 Assumptions and Approximations
This assumption only holds when $\langle dx/dt\rangle=0$ is possible, where $\langle\circ\rangle$ denotes the average of a quantity $\circ$ over time during a drug treatment cycle. Since the number of cells always increase during the off-drug time period, we need to have during the on-drug the following inequality:
@@
The right hand side is a finite positive value very close to 0, which is also increasing with $\epsilon$. This means that the tighter the control (i.e. the smaller the control window), the more difficult it becomes to maintain stabilization Eq. (30), as it reduces the allowable range for $x_{S}$. Therefore, when the adaptive treatment starts failing to suppress the cancer growth (i.e. the inequality does not hold anymore), we can make the following approximation: